# Layer-Wise Quantization: A Pragmatic and Effective Method for Quantizing LLMs Beyond Integer Bit-Levels

### v2

@@ Abstract
We present a simple variable quantization approach that quantizes different layers of a large language model (LLM) at different bit levels. Specifically, we quantize the most important layers to higher bit precision and less important layers to lower bits to achieve floating point quantization levels. We propose two effective strategies to measure the importance of layers within LLMs: the first measures the importance of a layer based on how different its output embeddings are from the input embeddings (the higher the better); the second estimates the importance of a layer using the number of layer weights that are much larger than average (the smaller the better). We show that quantizing different layers at varying bits according to our importance scores results in minimal performance drop with a far more compressed model size. Finally, we present several practical key takeaways from our variable layer-wise quantization experiments: (a) LLM performance under variable quantization remains close to the original model until 25-50% of layers are moved in lower quantization using our proposed ordering but only until 5-10% if moved using no specific ordering; (b) Quantizing LLMs to lower bits performs substantially better than pruning unless extreme quantization (2-bit) is used; and (c) Layer-wise quantization to lower bits works better in the case of larger LLMs with more layers compared to smaller LLMs with fewer layers. The code used to run the experiments is available at: https://github.com/RazvanDu/LayerwiseQuant.
@@
Layer-Wise Quantization: A Pragmatic and Effective Method for Quantizing LLMs Beyond Integer Bit-Levels
Razvan-Gabriel Dumitru1,2, Vikas Yadav1, Rishabh Maheshwary1, Paul-Ioan Clotan3, Sathwik Tejaswi Madhusudhan1, Mihai Surdeanu2 1ServiceNow Research, 2University of Arizona, 3Università di Bologna Correspondence: razvandumm@gmail.com
@@ 1 Introduction
Figure 1: Overall intuition behind our approach. We first rank the layers in an LLM (e.g., LLaMa-2-13B) using an importance score (shown here is ranking based on our Layer Input Modification (LIM) score, see section 3.1). The color intensity in each layer represent their LIM importance score. Then, the 30 most important layers are quantized in 4 bit while the remaining 10 least important layers are quantized in 2 bits, resulting in 3.5 bits as the average bit size.
@@
Large Language Models (LLMs) have achieved remarkable performance on a variety of tasks, especially when scaled to billions of parameters Jiang et al. (2023, 2024); Touvron et al. (2023); Zhang et al. (2022); Team et al. (2023). The largest open-source models available can have an upwards of 400B parameters, such as LLaMa3-400B Touvron et al. (2023). Even small models such as LLaMa3-8B require as much as 20GB of VRAM to run on a GPU at the original precision, making them unusable for low resource settings. In such settings model compression techniques such as quantization are critical Zhu et al. (2023b); Wan et al. (2023).
@@
Most prominent techniques for model compression broadly cover pruning Ma et al. (2023), knowledge distillation Gu et al. (2023), and quantization Zhu et al. (2023b). Pruning yields improvements in inference speed, but often results in substantial performance drop Frantar and Alistarh (2023); Men et al. (2024). On the other hand, quantization has proven to be a more robust solution for model size compression with comparatively much smaller performance drops Lin et al. (2024). In our work, we primarily focus on memory reduction through quantization. Further, quantization can be training specific or post-training Yao et al. (2024), where trained models are quantized without requiring any further training. We focus on post-training quantization due to its practicality.
@@
The majority of quantization techniques proposed recently (Frantar et al., 2023; Xiao et al., 2023; Yao et al., 2022, inter alia) focus on the quantization of all the LLM layers to a single precision bit, or quantizing only specific weights of the network Lin et al. (2024). This has shown to be costly, their effectiveness is data dependent, and their implementations can be considered moderately challenging. In contrast, we propose a simple technique that quantizes different layers at different bit precision depending on their importance. Figure 1 shows an example of our idea. We show that our approach is simple to implement, performs well, and provides greater flexibility to compress models in varying bit precision as per memory requirements. Further, our approach is complementary to other quantization techniques. For example, in our experiments we coupled our idea with two effective and prominent quantization techniques: GPT-Q and Quanto, but any other techinques can be used.
@@
Figure 2: Plots showing the effect of variable quantization for LLaMa2-13b and multiple datasets using Quanto. The leftmost point indicates LLM performance when all 40 layers of the LLM are represented in 4-bit; the rightmost point shows LLM performance when all layers are quantized to 2-bit. The dots on each curve (in each plot) show accuracy when the model is quantized to lower bits by converting less important layers to 2 bits one by one. Red and purple line indicate performance from 8bit and fp16 precision model (ceiling models). As shown, there is no considerable performance drop from fp16 or 8bit to 4bit precision. Hence, we focus our experiments on quantization below 4 bits. The vertical gray line indicates the quantization point that preserves 90% of the 4bit performance. The red line represents when layers are ordered randomly. We chose 3 random orders of the layers and quantized layers to 2 bits as per these orders. The standard deviation in performance from random orders are highlighted on the red curve. The curves are plotted on 2K evaluation data while results on full data is summarized in table 1.
Figure 3: Plots showing the effect of variable quantization for Mistral-7b and multiple datasets using Quanto. All notations are the same as in Figure 2.
@@
  • We introduce a novel method for layer-wise quantization of LLMs at different bit levels to achieve flexible lower bit precision overall, enabling LLMs to fit into range of specified memory capacity. We accompany our method with a detailed study, which highlights various important findings such as importance of layer ranking for quantizing less important layers with lower bits and more important layers with higher bit precision. We show that models can be quantized significantly more from 4-bit while retaining 90% of performance when following our proposed order.
@@
  • We propose and study two layer importance scores for variable quantization. Our first score named layer input modification (LIM) is based on how much a layer changes its input representations into the output ones. This score is calculated with an unlabeled text calibration corpus. The second scoring method, called z-score distribution (ZD), measures the distribution of weight magnitudes within a layer to determine its importance. To our knowledge, we are the first to propose layer orderings for quantization. We validate these scores by empirically showing that when LLM layers are quantized to lower bits as per rankings from our two importance score, they retain performance much more strongly than several other layer ordering baselines.
@@
  • We evaluate the impact of our variable quantization method based on layer importance on five top performing LLMs from different model families and different sizes. We draw several important practical findings from these experiments: (a) LLM performance under variable quantization remains close to the original model when using our ordering until reaching the level of 3.0-3.25 bits on average; (b) Quantizing LLMs to lower bits performs substantially better than pruning; however, under extreme quantization settings (i.e., $<=2-bits$) pruning shows slightly better results; (c) We show that quantization at two levels (i.e., quantizing some layers in $x$ bits and remaining layers in $y$ bits) performs much better than three levels of quantization, suggesting that the interaction between layers with different quantization is complex and may need more investigation; and (d) Layer-wise quantization to lower bits works better in the case of larger LLMs with more layers compared to smaller LLMs with fewer layers.
@@ 2 Related Work
Quantization techniques for large language models (LLMs) aim to reduce the precision of weights and activations to lower-bits without significantly compromising performance Zhu et al. (2023b). Popular approaches include Post-Training Quantization (PTQ) Banner et al. (2019) and Quantization-Aware Training (QAT) Liu et al. (2023). PTQ can be divided into static quantization, which uses a small dataset to calibrate scaling factors for weights and activations, and dynamic quantization, which quantizes activations on-the-fly during inference. Our study utilizes the static PTQ techniques such as GPT-Q in our experiments for practicality.
A few works have highlighted 4 bit precision as a robust quantization limit for wide variety of NLP tasks Dettmers and Zettlemoyer (2023). Our results (fig. 2 and 3) also show similar findings that performance drop is very minute from bf16 to 8 bits, and then to 4 bits. Hence, our experiments focus more on quantizing LLMs below 4 bits to highlight effectiveness of layer-wise quantization based on layer importance. Specifically, our empirical results (section 5) show that variable layer-wise quantization can retain 90% of the performance with a notable compression upto 2.85 bits overall. Concurrent (ArXiv preprint) work by Tai et al. (2024) also show effectiveness of variable quantization in vision language models.
A few previous works have studied layer importance in transformers Vaswani et al. (2017); Simoulin and Crabbé (2021). Some of the recent, (unpublished) concurrent works such as ShortGPT Men et al. (2024) have also proposed utilizing layer importance but primarily focusing on pruning. Shen et al. (2020) had utilized hessian information from each layer as an importance measure to quantize specific weight matrices at different bits. Different from these, our work is focused more on utilizing layer importance for quantizing more important layers (fully) in higher bits and less important layers in lower bits. Importantly, our proposed approach of layer-wise quantization of LLMs based on their layer importance is a method that can be easily extended with any quantization techniques like GPT-Q and Quanto (section 3.2). We also compare ShortGPT layer importance based pruning baseline with our layer-wise quantization approach in section 6.1.
@@ 3 Method
In our study, given a LLM with N layers denoted as $\{\mathbf{L_{1},L_{2},L_{3},....L_{N}}\}$, we first compute the importance of each layer, and then quantize them differently based on their importance. To achieve the former, we propose two scoring methods that allow us to rank the layers based on their respective importance scores.
@@ 3.1 Layer Importance Scores
We call our first layer importance score layer input modification (LIM). The intuition behind LIM is that the more a layer changes its received input embeddings the more important it must be. More formally, the LIM score for a specific layer $\mathbf{L_{i}}$ measures the modification of the representations of $\mathbf{L_{i}}$’s input (denoted by $\mathbf{L_{i}^{I}}$) to create its output representations (denoted by $\mathbf{L_{i}^{O}}$). A near similar score has been also studied for pruning (layer removal) in a concurrent (unpublished) work Men et al. (2024). The LIM score is the negative of cosine similarity between $\mathbf{L_{i}^{I}}$ and $\mathbf{L_{i}^{O}}$ as shown below:
@@
The dot product $\mathbf{L_{i}^{I}}\cdot\mathbf{L_{i}^{O}}$ quantifies the alignment between the input and output vectors, while the normalization factor $\|\mathbf{L_{i}^{I}}\|\|\mathbf{L_{i}^{O}}\|$ scales the dot product to the range of [-1, 1], corresponding to the cosine of the angle between the two vectors. The negative sign in the LIM score indicates that a higher similarity (cosine similarity closer to 1) between $\mathbf{L_{i}^{I}}$ and $\mathbf{L_{i}^{O}}$ results in a lower importance score. For measuring change in $\mathbf{L_{i}^{I}}$ to $\mathbf{L_{i}^{O}}$, we use all 50 documents of pg19! Rae et al. (2019) (an unlabelled text corpus) as calibration corpus.
@@
Layer ranking baselines:
To highlight the benefits of our layer importance scores, we compare them to two baselines. First, we randomly rank layers with different random seeds to highlight ordering of layer is important. Second, to specifically highlight the strengths of the LIM score, we consider a reverse LIM ordering as a baseline. Finally, we also considered the reverse index of layers, i.e., from layer 30 to layer 1, following Gromov et al. (2024) who has shown just removing layers from the top of the LLM to be an effective order for layer pruning. We implement the same pruning with specified reverse indexes from Gromov et al. (2024) and show its comparison in Figure 4(a).
@@ 3.2 Quantization Techniques
In our work, we employ two post-training quantization techniques—GPT-Q and Quanto—to our proposed layer-wise quantization study.
@@ 3.2.1 Quanto
Quanto is a fast-acting quantization method [^1] https://github.com/huggingface/optimum-quanto that simplifies the process of reducing precision after training. In practice, Quanto achieves quicker quantization by applying uniform scaling factors across all layers of a model, avoiding the need for detailed data-driven analysis of each layer’s distribution. Similarly to most techniques, it allows for int2, int4, int8, and fp8 quantization. In our implementation for variable quantization of different layers, we quantized the same model to two different bit levels (lets say int2 and int4). Then, based on the layer importance order, we selected less important layers from the int2 quantized layer sets and more important layers from int4 quantized sets.
@@
Table 1: Accuracy on full evaluation datasets of different models quantized with Quanto. All layers start at 4 bits; we then quantize N number of layers in 2 bits where N is mentioned in the “Layers 2 bits” column. We also show results for 8 to 4 bit quantization in Appendix in table 4. Average performances within 90% of the 4 bit model are highlighted in bold.
@@
Table 2: Comparison of different models and their performance across various tasks with GPT-Q quantization.

### v3

@@ Abstract
We present a simple meta quantization approach that quantizes different layers of a large language model (LLM) at different bit levels, and is independent of the underlying quantization technique. Specifically, we quantize the most important layers to higher bit precision and less important layers to lower bits. We propose two effective strategies to measure the importance of layers within LLMs: the first measures the importance of a layer based on how different its output embeddings are from the input embeddings (higher is better); the second estimates the importance of a layer using the number of layer weights that are much larger than average (smaller is better). We show that quantizing different layers at varying bits according to our importance scores results in minimal performance drop with a far more compressed model size. Finally, we present several practical key takeaways from our variable layer-wise quantization experiments: (a) LLM performance under variable quantization remains close to the original model until 25–50% of layers are moved in lower quantization using our proposed ordering but only until 5–10% if moved using no specific ordering; (b) Adding layer importance to inherently dynamic quantization techniques can further improve their performance, showing that our approach is complementary to other dynamic quantization methods; (c) Quantizing LLMs to lower bits performs substantially better than pruning unless extreme quantization (2-bit) is used; and (d) Layer-wise quantization to lower bits works better in the case of larger LLMs with more layers compared to smaller LLMs with fewer layers. Our code is publicly available at https://github.com/RazvanDu/LayerwiseQuant/.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance on a variety of tasks, especially when scaled to billions of parameters Jiang et al. (2023; 2024); Touvron et al. (2023); Zhang et al. (2022); Team et al. (2023). The largest open-source models available can have an upwards of 400B parameters, such as LLaMa3-400B Touvron et al. (2023). Even small models such as LLaMa3-8B require as much as 20GB of VRAM to run on a GPU at the original precision, making them unusable for low resource settings. In such settings model compression techniques such as quantization are critical Zhu et al. (2023); Wan et al. (2023).
@@
Most prominent techniques for model compression broadly cover pruning Ma et al. (2023), knowledge distillation Gu et al. (2023), and quantization Zhu et al. (2023). Pruning yields improvements in inference speed, but often results in substantial performance drop Frantar & Alistarh (2023); Men et al. (2024). On the other hand, quantization has proven to be a more robust solution for model size compression with comparatively much smaller performance drops Lin et al. (2024). In our work, we primarily focus on memory reduction through quantization. Further, quantization can be training specific or post-training Yao et al. (2024), where trained models are quantized without requiring any further training. We focus on post-training quantization due to its practicality.
@@
The majority of quantization techniques proposed recently (Frantar et al., 2023; Xiao et al., 2023; Yao et al., 2022, inter alia) focus on the quantization of all the LLM layers to a single precision bit, or quantizing only specific weights of the network Lin et al. (2024). This has shown to be costly, their effectiveness is data dependent, and their implementations can be considered moderately challenging. In contrast, we propose a simple meta quantization technique that quantizes different layers at different bit precision depending on their importance. Figure 5 shows an example of our idea. We show that our approach is simple to implement, is independent of the underlying quantization technique (e.g., we experimented with two prominent quantization techniques: GPT-Q (Frantar et al., 2023) and Quanto [^1] https://github.com/huggingface/optimum-quanto), performs well, and provides greater flexibility to compress models in varying bit precision as per memory requirements.
Figure 1: Overall intuition behind our approach. We first rank the layers in an LLM (e.g., LLaMa-2-13B here) in descending order of an importance score (shown here is ranking based on our Layer Input Modification (LIM) score, see section 3.1). The color intensity of each layer, which represents their LIM importance score (darker color indicates higher importance score), highlights that the original layer structure (left-hand side of the figure) does not have the layers sorted according to their importance. This observation holds for several other LLMs (see Figure 6 in the appendix). After sorting (right-hand side of the figure), the 30 most important layers are quantized in 4 bits while the remaining 10 least important layers are quantized in 2 bits, resulting in 3.5 bits as the average bit size.
@@
  • We introduce a novel method for layer-wise quantization of LLMs at different bit levels to achieve flexible lower bit precision overall. Our method is designed to maximize performance under a given memory budget by keeping as many layers as possible in higher precision. We accompany our method with a detailed study, which highlights various important findings such as importance of layer ranking for quantizing less important layers with lower bits and more important layers with higher bit precision. We show that models can be quantized significantly more from 4-bit while retaining 90% of performance when following our proposed order.
@@
  • We propose and study two layer importance scores for variable quantization. To our knowledge, we are the first to propose layer orderings for quantization. Our first score, named layer input modification (LIM), is based on how much a layer changes its input representations into the output ones. This score is calculated with an unlabeled text calibration corpus. The second scoring method, called z-score distribution (ZD), measures the distribution of parameter weights within a layer to determine its importance. Thus, ZD does not require calibration data. We validate these scores by empirically showing that when LLM layers are quantized to lower bits as per rankings from our two importance scores, they retain performance much more strongly than several other layer ordering baselines. We observe that LIM performs better than ZD on average but the differences are not large, and ZD has less performance variance for different LLMs. We consider this a success for ZD, which is simpler and does not require calibration data.
@@
  • We evaluate the impact of our variable quantization method based on layer importance on five top performing LLMs from different model families and different sizes. We draw several important practical findings from these experiments: (a) LLM performance under variable quantization remains close to the original model when using our ordering until reaching the level of 3.0–3.25 bits on average; (b) Quantizing LLMs to lower bits performs substantially better than pruning; however, under extreme quantization settings (i.e., <=<=< = 2-bits) pruning shows slightly better results; (c) We show that quantization at two levels (i.e., quantizing some layers in $x$ bits and remaining layers in $y$ bits) performs much better than three levels of quantization, suggesting that the interaction between layers with different quantization is complex and may need more investigation; and (d) Layer-wise quantization to lower bits works better in the case of larger LLMs with more layers compared to smaller LLMs with fewer layers.
Figure 2: Plots showing the effect of variable quantization for LLaMa2-13b and multiple datasets using Quanto. The leftmost point indicates LLM performance when all 40 layers of the LLM are represented in 4-bit; the rightmost point shows LLM performance when all layers are quantized to 2-bit. The dots on each curve (in each plot) show accuracy when the model is quantized to lower bits by converting less important layers to 2 bits one by one. Red and purple line indicate performance from 8bit and fp16 precision model (ceiling models). As shown, there is no considerable performance drop from fp16 or 8-bit to 4-bit precision. Hence, we focus our experiments on quantizing below 4 bits. The vertical gray line indicates the quantization point that preserves 90% of the 4-bit performance. The red line represents when layers are ordered randomly. We chose 3 random orders of the layers and quantized layers to 2 bits as per these orders. The standard deviation in performance from random orders are highlighted on the red curve. The curves are plotted on 2K evaluation data while results on full data is summarized in table 1. The figure shows that our method retains performance much better under more aggressive quantization than all baselines.
Figure 3: Plots showing the effect of variable quantization for Mistral-7b and multiple datasets using Quanto. All notations are the same as in Figure 2. Again, the figure shows that our method retains performance much better under more aggressive quantization than all baselines.
@@ 2 Related Work
Quantization techniques for large language models (LLMs) aim to reduce the precision of weights and activations to lower-bits without significantly compromising performance (Zhu et al., 2023). Popular approaches include Post-Training Quantization (PTQ) (Banner et al., 2019) and Quantization-Aware Training (QAT) (Liu et al., 2023). PTQ can be divided into static quantization, which uses a small dataset to calibrate scaling factors for weights and activations, and dynamic quantization, which quantizes activations on-the-fly during inference. Our study utilizes the static PTQ techniques such as GPT-Q (Frantar et al., 2023) in our experiments for practicality, but, importantly, it is agnostic to the actual quantization technique used.
A few works have highlighted 4-bit precision as a robust quantization limit for wide variety of NLP tasks (Dettmers & Zettlemoyer, 2023). Our results (fig. 2 and 3) also show similar findings that performance drop is small from bf16 to 8-bits, and then to 4-bits. Hence, our experiments focus more on quantizing LLMs below 4-bits to highlight the effectiveness of layer-wise quantization based on layer importance. Specifically, our empirical results (section 5) show that variable layer-wise quantization can retain 90% of the performance with a notable compression up to 2.85-bits overall. Concurrent (ArXiv preprint) work by Tai et al. (2024) also show effectiveness of variable quantization in vision language models.
A few previous works have studied layer importance in transformers (Vaswani et al., 2017; Simoulin & Crabbé, 2021). Some of the recent, (unpublished) concurrent works such as ShortGPT (Men et al., 2024) have also proposed utilizing layer importance but primarily focusing on pruning. Shen et al. (2020) had utilized hessian information from each layer as an importance measure to quantize specific weight matrices at different bits. Different from these, our dynamic strategy focuses on utilizing layer importance for quantizing more important layers in higher bits and less important layers in lower bits. Importantly, our proposed approach of layer-wise quantization of LLMs based on their layer importance is a meta method that can be coupled with any quantization techniques such as GPT-Q and Quanto (section 3.3). We also compare ShortGPT’s layer importance-based pruning with our layerwise quantization approach in section 6.1, and show that they are largely complementary.
@@ 3 Method
At a high level, our meta quantization strategy operates in two steps: first, given an LLM with N layers denoted as $\{\mathbf{L_{1},L_{2},L_{3},....L_{N}}\}$, we first compute the importance of each layer. Second, we then employ an existing quantization technique to quantize layers differently based on their importance.
@@ 3.1 Layer Importance Scores
We propose two layer importance scoring methods.
We call our first layer importance score layer input modification (LIM). The intuition behind LIM is that the more a layer changes its received input embeddings the more important it must be. More formally, the LIM score for a specific layer $\mathbf{L_{i}}$ measures the modification of the representations of $\mathbf{L_{i}}$’s input (denoted by $\mathbf{L_{i}^{I}}$) to create its output representations (denoted by $\mathbf{L_{i}^{O}}$). A near similar score has been also studied for pruning (layer removal) in a concurrent (unpublished) work (Men et al., 2024). The LIM score is the negative of cosine similarity between $\mathbf{L_{i}^{I}}$ and $\mathbf{L_{i}^{O}}$ as shown below:
@@
The dot product $\mathbf{L_{i}^{I}}\cdot\mathbf{L_{i}^{O}}$ quantifies the alignment between the input and output vectors, while the normalization factor $\|\mathbf{L_{i}^{I}}\|\|\mathbf{L_{i}^{O}}\|$ scales the dot product to the range of [-1, 1], corresponding to the cosine of the angle between the two vectors. The negative sign in the LIM score indicates that a higher similarity (cosine similarity closer to 1) between $\mathbf{L_{i}^{I}}$ and $\mathbf{L_{i}^{O}}$ results in a lower importance score. For measuring change in $\mathbf{L_{i}^{I}}$ to $\mathbf{L_{i}^{O}}$, we use all 50 documents of pg19! Rae et al. (2019a) (an unlabelled text corpus) as calibration corpus.
@@
Note that both these two scores have distinct advantages and disadvantages. LIM relies on runtime information (i.e., changes in representations) so it is more aligned with inference behavior. However, LIM requires a tuning dataset to compute these representations. In contrast, ZD does not require calibration data as it uses solely the network parameters.
Baselines:
To highlight the benefits of our layer importance scores, we compare them to three baselines. The first baseline simply randomly ranks layers. [^2] We use multiple different random seeds for stability. The second baseline implements a reverse LIM ordering. Finally, to compare our approach with pruning, we implemented the pruning strategy of Gromov et al. (2024), who has shown that removing layers from the top of the LLM to be an effective pruning strategy.
3.2 Choosing the Number of Layers in Higher Precision Given a Memory Budget
Under a fixed memory budget, our approach aims to keep as many layers as possible in higher precision. The maximum number of layers to keep in higher precision given available memory is easy to compute. In particular, assume we have $M_{available}$ memory available. Having all $N_{layers}$ layers in the network in lower precision requires $M_{lower}$ memory, while having all of them in higher precision requires $M_{higher}$ memory. The maximum number of layers that can be kept in higher precision is then calculated as:
$N_{higher}=\lfloor\frac{M_{available}-M_{lower}}{M_{higher}-M_{lower}}*N_{layers}\rfloor$
For example, if we have 20GB of VRAM available; a model with 32 layers takes 17GB in lower precision (e.g., 2-bits) and 34GB in higher precision (e.g., 4-bits), then the number of layers using higher precision is: $N_{higher}=\lfloor\frac{20-17}{34-17}*32\rfloor=5$. This means that 5 of the layers can be in 4-bit precision and the other $32-5=27$ have to be in 2-bit precision, for a total memory of $19.65GB<20GB$. This also underscores the practicality of our technique, as we do not add an extra hyper-parameter that requires tuning and we make use of most of the available memory.
@@ 3.3 Quantization Techniques
To show that our meta quantization approach is independent of the underlying quantization technique, we couple it with two well-known post-training quantization techniques: GPT-Q and Quanto. While our focus is not on low-level quantization methods, we summarize both methods below for completeness.
@@ 3.3.1 Quanto
Quanto is a fast-acting quantization method [^3] https://github.com/huggingface/optimum-quanto that simplifies the process of reducing precision after training. In practice, Quanto achieves quicker quantization by applying uniform scaling factors across all layers of a model, avoiding the need for detailed data-driven analysis of each layer’s distribution. Similarly to most techniques, it allows for int2, int4, int8, and fp8 quantization. The Quanto library is part of Hugging Face, with symmetric per-tensor or per-channel projection for int8 and float8, and group-wise affine (with a shift or ’zero-point’) for lower bit-widths, such as int2 or int4. In this paper, we used Quanto in the 2-bit and 4-bit setting, utilizing an asymmetric quantization method with RTN (round to the nearest), with the exception of Table 5 where we used 4-bit and 8-bit settings. In our implementation for variable quantization of different layers, we quantized the same model to two different bit levels (int2 and int4). Then, based on the layer importance order, we selected less important layers from the int2 quantized sets and more important layers from the int4 quantized sets.
@@
Table 1: Accuracy on full evaluation datasets of different models quantized with Quanto. All layers start at 4-bits; we then quantize N number of layers in 2-bits where N is mentioned in the “Layers 2-bits” column. We also show results for 8 to 4 bit quantization in Appendix in table 5. Average performances within 90% of the 4-bit model are highlighted in bold.
@@
Table 2: Comparison of 4 models and their performance across various tasks with GPT-Q quantization.